OpenStudy (anonymous):

The length of a rectangle is 4 more than the width. The area is 60 square inches. Find the dimensions

OpenStudy (anonymous):

60 = (w + 4)w 0 = w^2 + 4w - 60 Then you solve that for width, and then add 4 to get the length. PROTIP: Don't use the negative root for w.

OpenStudy (anonymous):

i got 0=(w+10)(w-6)

OpenStudy (anonymous):

so when i take 0=w+10 that means w=-10?

OpenStudy (anonymous):

and the other is 6?

OpenStudy (anonymous):

i know the width is 6 but the length should be 24 to make it all equal 60 right>

OpenStudy (anonymous):

Yeah. So the width is 6. And the length is 10. 6 * 10 = 60

OpenStudy (anonymous):

ok thanks!